OpenStudy (oleg3321):

Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown

OpenStudy (oleg3321):

theres a graph -2,0 and 0,6

OpenStudy (oleg3321):

y =1/3 x y = -1/3 x + 4 y = 3x + 20 y = -3x - 16

OpenStudy (oleg3321):

@Astrophysics

OpenStudy (oleg3321):

@mathstudent55

OpenStudy (oleg3321):

@robtobey

OpenStudy (oleg3321):

@No.name

OpenStudy (oleg3321):

@ganeshie8

OpenStudy (oleg3321):

please help will give medal

OpenStudy (driftracer305):

k perpendicular line has a slope that is oppo. reciprocal of the line given

OpenStudy (oleg3321):

thats the graph

OpenStudy (oleg3321):

@ganeshie8

OpenStudy (oleg3321):

please give answer

OpenStudy (driftracer305):

B dude...

OpenStudy (oleg3321):

are you sure

OpenStudy (driftracer305):

500%

OpenStudy (oleg3321):

okay heres your medal

OpenStudy (oleg3321):

thanks

OpenStudy (driftracer305):

uh pleasure